OpenStudy (anonymous):

Determine how many slices you will cut your pizza into. You need to pick two different numbers of slices. Once you have determined the number of slices, calculate the following for both ways of cutting your pizza: • the interior angle and arc length of the slices • the area of each slice

OpenStudy (anonymous):

Determine the delivery radius for your shop. Draw a point on a coordinate plane where your shop will be located. Create two different radii lengths from your shop, and construct the circles that represent each delivery area. How much area will each delivery radius cover? Write the equation for each circle

hero (hero):

Seems like one of those "do it yourself" problems to me.

OpenStudy (anonymous):

I really don't understand how to do this tho :(

hero (hero):

@inkyvoyd, help her

OpenStudy (inkyvoyd):

Let's start with a couple of things 1. The formula for the area of a circle and sector 2. The formula for the arc length of a circle 3. The equation of a circle centered at (h,k) @Hero , your turn. Alternatively, @lgbasallote

hero (hero):

I'm too busy. I don't have any time.

OpenStudy (lgbasallote):

@Hero is adequate enough though

hero (hero):

Sorry, maybe on the weekends

OpenStudy (inkyvoyd):

Stop being lazy guys (I'm not lazy :P) @lgbasallote , helpy out.

OpenStudy (anonymous):

Just please somebody help :( I've been stuck on this problem for the longest time.

OpenStudy (inkyvoyd):

I got like 9 minutes, and I have about 3 days worth of online coursework to catch up on for the moment. Let me retag that igbiw. @lgbasallote

OpenStudy (inkyvoyd):

@dpaInc

OpenStudy (inkyvoyd):

@nbouscal

OpenStudy (inkyvoyd):

@ParthKohli . My last chance. Help out please

OpenStudy (zepp):

First, pick any random number to determine the size of the pizza, we know that a circle is defined by 1 thing, the radius; then, pick the number of slices.

OpenStudy (zepp):

Let me make the first example, I have a 14cm radius pizza, and I want 3 slices. First off, we know that a circle has 360 degrees inside. 3 equal slices, so 360/3 = 120 degrees for each slice.  Then, to find a arc length of each slice, we can use the little formula we are going to invent by ourselves, Circonference \(\div\) Number of slice = Arc length of each slice The formula to find the circonference of a circle is \(2 \pi r\) Let's plug our radius in, \(2 \pi 14=28 \pi\) Okay, cool, we have our circonference, 3 slices. Apply our little formula, Circonference \(\div\) Number of slice = Arc length of each slice; \(28 \pi \div 3 \approx 29.32cm\) So the arc length of each slice would be 29.32cm. Next, the area, simple use the formula \(\pi r^2\).